Karen and steve start traveling toward each other from 825 miles apart. karen is traveling at 32 mph and steve at 36 miles per hour. compute the distance karen will travel befor they meet
Found 2 solutions by rfer, fcabanski:
Answer by rfer(16322)

approaching at 68mph
825/68=12.1323 hrs
12.1323*32=388.25 miles (Karen)

Answer by fcabanski(1391)

The time they travel to reach this point is the same, and D=rt is the formula for distance, rate and time. Divide both sides by r to find t=D/r


In K's case she will travel some distance, x (D=x, a variable we don't know), so the t = x/32.


In Steve's case he'll travel 825-x (the full distance, minus whatever K travels, which is where they will meet. For example if Karen travels 100 miles, then Steve must travel 825-100 = 725 miles to meet her.) The same t, the time they travel to meet, is t= (825-x)/36.


Those two times are equal, so set the equations = to each other and solve for x.


x/32 = (825-x)/36 Cross Multiply 36x = 26400 - 32x


Bring the x's to the left...68x = 26400 Divide by 68 and x=388.24.


Karen travels 388.24 miles before they meet.